# Sequential model confidence sets

### v1

@@ 3.2 SMCSs for the strongly superior models
$E_{i\cdot,t}=\frac{1}{m-1}\sum_{j\neq i}E_{ij,t},\quad t\in\mathbb{N},$
which is also an e-process for $\mathcal{H}^{\textrm{s}}_{i\cdot}$ as the arithmetic mean of $m-1$ e-processes. Finally, we adjust the e-processes
$(E_{1\cdot t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}},$ (9)
symmetrically by the closure principle in the spirit of Vovk and Wang (2021), as explained in detail in Appendix C, to obtain family-wise valid e-processes $(E_{1\cdot,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t}^{\star})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}^{\textrm{s}}_{1\cdot},\dots,\mathcal{H}^{\textrm{s}}_{m\cdot}$. We use the arithmetic mean as the $e$-merging function as it essentially dominates all symmetric e-merging functions (Vovk and Wang, 2021, Proposition 3.1.). For $\alpha\in(0,1)$, we apply the sequential tests given at (8) and let
@@
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}^{\star}\leq 1/\alpha\right\}.$ (10)
@@
That is, at each $t\in\mathbb{N}$, we exclude model $i\in\mathcal{M}_{0}$ from the SMCS if we may reject the hypothesis $\mathcal{H}^{\textrm{s}}_{i\cdot}$ according to the sequential testing procedure, and include it otherwise.
Lemma 3.1.
@@
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (10) is a sequence of model confidence sets at level $\alpha$ for the sequence of strongly superior objects $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\in\mathbb{N}}$, and so is its running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$.
@@
For any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, we have
$\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\textrm{s},\star}_{t}\not\subseteq\widehat{\mathcal{M}}_{t})=\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{i\cdot,t}^{\star}>1/\alpha\textrm{ for some }i\in\mathcal{M}^{\textrm{s},\star}_{t}\right).$
Moreover, $i\in\mathcal{M}^{\textrm{s},\star}_{t}$ if and only if $\mathbb{Q}\in\mathcal{H}^{\textrm{s}}_{i\cdot,t}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,r}\leq 0,\forall r\leq t,\forall j\in\mathcal{M}_{0}\}$. Since $(E_{1\cdot,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t}^{\star})_{t\in\mathbb{N}}$ are FWV, there exists a dominating family of e-processes $\{(E_{t}^{\mathbb{Q}})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$ such that $E_{r}^{\mathbb{Q}}\geq E_{i\cdot,r}^{\star}$ for all $r\leq t$ and all $i\in\mathcal{M}^{\textrm{s},\star}_{t}$. So, we may conclude
$\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\textrm{s},\star}_{t}\not\subseteq\widehat{\mathcal{M}}_{t})$ $=\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{i\cdot,t}^{\star}>1/\alpha\textrm{ for some }i\in\mathcal{M}^{\textrm{s},\star}_{t}\right)$
$\leq\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{t}^{\mathbb{Q}}>1/\alpha\right)\leq\alpha.$
It follows directly, that the running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$, is a SMCS as well, since $i\notin\mathcal{M}^{\star}_{t_{0}}$ implies $i\notin\mathcal{M}^{\star}_{t}$ for all $t\geq t_{0}.$

### v3

@@ 3.2 SMCSs for the strongly superior and uniformly weakly superior objects
$E^{\star}_{i\cdot,t}=\min_{I\subseteq\{1,\dots,m\}:i\in I}\frac{1}{\lvert I\rvert}\sum_{j\in I}E_{j\cdot,t}\leq E_{i\cdot,t},\quad i\in\mathcal{M}_{0},t\in\mathbb{N},$ (6)
see Supplement C for further details. As shown in Vovk and Wang (2021, Proposition 3.1.), the arithmetic mean essentially dominates all other symmetric e-merging functions. Algorithm 1 of Vovk and Wang (2021) allows for an efficient computation of the minimum at (6). Finally, for some significance level $\alpha\in(0,1)$, we define
@@
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}^{\star}<1/\alpha\right\},\quad t\in\mathbb{N}.$ (7)
@@
In words, at each $t\in\mathbb{N}$, we exclude model $i\in\mathcal{M}_{0}$ if we may reject the hypothesis $\mathcal{H}^{\bullet}_{i\cdot}$ according to the sequential test $\psi_{i\cdot,t}^{\star}=\mathbbm{1}\{E_{i\cdot,t}^{\star}\geq 1/\alpha\}$, and include it otherwise. The proof the the following theorem is given in Supplement C.
Theorem 3.1.
@@
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (7) is an SMCS at level $\alpha$ for $\mathcal{M}^{\bullet,\star}$, $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, and so is its running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$.
Theorem 3.1 works since we control the family-wise error rate of the tests $\psi_{i\cdot,t}^{\star}$ over all models $i$. Family-wise error rate control relies on the closure principle (Marcus et al., 1976) and some e-merging function, see Supplement C. Alternatively, one could adjust the pairwise e-processes $E_{ij,t}$ directly to obtain adjusted e-processes for the pairwise hypotheses $\mathcal{H}^{\bullet}_{ij}$ and exclude the model $i$ from the SMCS at $t\in\mathbb{N}$ if $E^{\star}_{ij,t}\geq 1/\alpha$ for some $j\neq i$. However, the corresponding tests can be shown to be uniformly less powerful, see Supplement C.
Remark 1.
Except from the family-wise error rate, the false discovery rate (FDR) is arguably the most commonly used criterion in multiple testing (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001). Recently, Wang and Ramdas (2022) studied FDR control for e-values. In principle, SMCSs could be constructed based on FDR control. However, this fundamentally changes the type of validity we are obtaining for ^$\widehat{\mathcal{M}}_{t}$. It allows us to bound the fraction of wrongly rejected models by the total number of rejected models at all time points, which is different from our target coverage guarantee at (4). We give some further details in Supplement D.
The following proposition provides e-processes for the strong hypotheses under the assumption of conditionally bounded loss differences. Together with Theorem 3.1, it directly leads to the derivation SMCSs for the strongly superior objects.
Proposition 3.2.
Assume that $|d_{ij,t}|\leq c_{ij,t}/2$ for $i,j\in\mathcal{M}_{0}$, $t\in\mathbb{N}$, and some predictable sequence $(c_{ij,t})_{t\in\mathbb{N}}\subseteq(0,\infty)$. Then, for any $i,j\in\mathcal{M}_{0}$, $E_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}d_{ij,r})$ is an e-process for $\mathcal{H}_{ij}^{\textrm{s}}$, for any predictable $(\lambda_{ij,t})_{t\in\mathbb{N}}$ with $0\leq\lambda_{ij,t}\leq 1/c_{ij,t}$.
@@
The process $M_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}(d_{ij,r}-\mu_{ij,r}))$ is a nonnegative martingale with $M_{ij,0}=1$. The claim follows by observing that, for $\mathbb{Q}\in\mathcal{H}^{\textrm{s}}_{ij}$, $M_{ij,t}\geq E_{ij,t}$ for all $t\in\mathbb{N}$. ∎
The following proposition is due to Choe and Ramdas (2023, Theorem 3), and allows to construct e-processes for the weak hypothesis under the assumption of uniformly bounded loss differences. Together with Theorem 3.1, it provides SMCSs for the uniformly weakly superior objects. In Section 3.4, we explain how we can weaken the assumption of uniformly bounded loss differences to conditionally bounded loss differences.
Proposition 3.3.
Let $|d_{ij,t}|\leq c_{ij}/2$ for $i,j\in\mathcal{M}_{0}$, $t\in\mathbb{N}$, and some $c_{ij}>0$. Then, for any $i,j\in\mathcal{M}_{0}$,
$E_{ij,t}=\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij,t}-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\},\quad t\in\mathbb{N},$
is an e-process for $\mathcal{H}_{ij}^{\textrm{uw}}$, for any $0\leq\lambda_{ij}\leq 1/c_{ij}$ and
$\psi_{E,c_{ij}}(\lambda)=(-\log(1-c_{ij}\lambda_{ij})-c_{ij}\lambda_{ij})/c_{ij}^{2}\quad\textrm{and}\quad V_{ij,t}=\sum_{r=1}^{t}(d_{ij,r}-\gamma_{ij,r})^{2},$ (8)
where $(\gamma_{ij,t})_{t\in\mathbb{N}}$ with $\lvert\gamma_{ij,t}\lvert\leq c_{ij}/2$ is a predictable sequence.
3.3 SMCSs for the weakly superior objects
Our SMCS construction for the weakly superior objects defined at (3) builds on time-uniform confidence regions. Proofs for this section are given in Supplement E.
For $i,j\in\mathcal{M}_{0}$, $t\in\mathbb{N}$, $x\in\mathbb{R}$, let $M_{ij,t}(x)$ be such that $(M_{ij,t}(\Delta_{ij,t}))_{t\in\mathbb{N}}$ is a test supermartingale for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. Let $\mathbb{R}_{0}^{m\times m}$ be the family of all $m\times m$-matrices with diagonal entries equal to zero and define for $\boldsymbol{X}=(x_{ij})_{ij}\in\mathbb{R}_{0}^{m\times m}$,
With $\boldsymbol{\Delta}_{t}=(\Delta_{ij,t})_{ij}\in\mathbb{R}_{0}^{m\times m}$, also $(M_{t}(\boldsymbol{\Delta}_{t}))_{t\in\mathbb{N}}$ is a test supermartingale for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. Therefore, for any $\alpha\in(0,1)$, $C_{t}=\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid M_{t}(\boldsymbol{X})\leq 1/\alpha\}$ is a $(1-\alpha)$-confidence sequence for $(\boldsymbol{\Delta}_{t})_{t\in\mathbb{N}}$. Indeed, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, by Ville’s inequality,
$\mathbb{Q}(\exists t\in\mathbb{N}:\boldsymbol{\Delta}_{t}\notin C_{t})=\mathbb{Q}(\exists t\in\mathbb{N}:M_{t}(\boldsymbol{\Delta}_{t})>1/\alpha)\leq\alpha.$
We construct an SMCS from $C_{t}$ as follows. If there exists a $j\not=i$ such that we may reject $\Delta_{ij,t}\leq 0$, we exclude $i$ from the SMCS. This is the case, if and only if there exists $j\not=i$ such that $C_{t}\cap\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid x_{ij}\leq 0\}=\emptyset$. Therefore, we define
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t}\cap\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid x_{ij}\leq 0\}\not=\emptyset\text{ for all }j\neq i\right\},\quad t\in\mathbb{N}.$ (9)
Theorem 3.4.
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (9) is an SMCS at level $\alpha$ for the weakly superior objects $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$. Its running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$ is an SMCS at level $\alpha$ for the uniformly weakly superior objects $\mathcal{M}^{\textrm{uw},\star}$.
For computation of the SMCS, it is useful to understand the shape of the sets $C_{t}$. The following result treats a relevant case that allows for simpler computations.
Proposition 3.5.
Suppose that the functions $x\mapsto M_{ij,t}(x)$ are nonnegative, convex and decreasing for all $i,j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$. Then, $C_{t}$ is a convex upper set, that is, for $\boldsymbol{A},\boldsymbol{B}\in C_{t}$, it follows that $\lambda\boldsymbol{A}+(1-\lambda)\boldsymbol{B}\in C_{t}$ for all $\lambda\in[0,1]$, and $\boldsymbol{A}\in C_{t}$ and $\boldsymbol{A}\leq\boldsymbol{C}$ implies that $\boldsymbol{C}\in C_{t}$, where the inequality is understood componentwise.
Under the conditions of Proposition 3.5, the definition of ^$\widehat{\mathcal{M}}_{t}$ at (9) simplifies:
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t}\cap\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid x_{ij}=0\}\not=\emptyset\text{for all }j\neq i\right\},\quad t\in\mathbb{N}.$ (10)
For uniformly bounded loss differences, the following proposition provides test supermartingales that satisfy the assumptions of Proposition 3.5. It can be found in Choe and Ramdas (2023, Proposition 1 of Appendix). In the next section, we discuss how to accommodate conditionally bounded loss differences.
Proposition 3.6.
Let $|d_{ij,t}|\leq c_{ij}/2$ for $i,j\in\mathcal{M}_{0}$, $t\in\mathbb{N}$, and some $c_{ij}>0$. Then, for any $i,j\in\mathcal{M}_{0}$, $(M_{ij,t}(\Delta_{ij,t}))_{t\in\mathbb{N}}$ is a test supermartingale for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, where
$M_{ij,t}(x)=\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij,t}-\lambda_{ij}tx-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\},\quad t\in\mathbb{N},$
for $0\leq\lambda_{ij}\leq 1/c_{ij}$ and $\psi_{E,c_{ij}}$ and $V_{ij,t}$ are given at (8).
Since $\mathcal{M}^{\text{uw},\star}=\bigcap_{t=1}^{\infty}\mathcal{M}_{t}^{\text{w},\star}$, one could alternatively use the running intersection of the SMCSs defined at (10) to construct SMCSs for the uniformly weakly superior objects. However, the SMCSs defined at (10) are much more expensive computationally. Therefore, we only compute SMCSs for uniformly weakly superior objects using Theorem 3.1.
3.4 Predictable bounds and betting schemes
Our methods for the (uniformly) weakly superior objects require that the score differences are uniformly bounded over time, which is clearly more restrictive than the assumption of conditionally bounded score differences imposed for the strongly superior objects. However, one can always transform conditionally bounded score differences into uniformly bounded ones thereby modifying the underlying loss and corresponding superior objects.
Assume that $\lvert d_{ij,t}\rvert\leq c_{ij,t}/2$ for all $t\in\mathbb{N}$ for some predictable $(c_{ij,t})_{t\in\mathbb{N}}$. Then, the transformed loss differences $\tilde{d}_{ij,t}=d_{ij,t}/c_{ij,t},t\in\mathbb{N}$ are uniformly bounded with constant $c=1$ that is $\lvert\tilde{d}_{ij,t}\rvert\leq 1/2$ for all $t\in\mathbb{N}$. However, if we use the transformed loss differences, we target (possibly) different sequences of superior objects ~$\tilde{\mathcal{M}}^{\textrm{uw},\star}=\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\text{for all}\;j\in\mathcal{M}_{0},t\in\mathbb{N}\}$ and ~$\tilde{\mathcal{M}}^{\textrm{w},\star}_{t}=\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\text{for all}\;j\in\mathcal{M}_{0}\}$ for $\tilde{\Delta}_{ij,t}=(1/t)\sum_{r=1}^{t}\mathbb{E}(\tilde{d}_{ij,r}\mid\mathcal{F}_{r-1})=(1/t)\sum_{r=1}^{t}\mu_{ij,r}/c_{ij,r}.$
In Section 5, we use the given transformation to convert conditionally bounded CRPS differences of wind gust forecasts into uniformly bounded loss differences. Importantly, for any proper scoring rule (consistent scoring function) ${\rm S}$, the scaled function $\tilde{{\rm S}}={\rm S}/c$ is proper (consistent) as well for any $c>0$, an observation which justifies the transformation from a theoretical perspective. This property can be extended in the sequential setting, see Supplement A. With the given scaling, all observations have the same maximal impact on the ranking, an effect which seems appealing for many (but surely not for all) applications.